A set of cards contains four suits. Two suits are​ red, and two suits are black. In each suit there are cards numbered from 2 to 10 as well as a​ jack, a​ queen, a​ king, and an ace. Consider the following probability for one card selected at random.

​P(the selected card is a three or a red ​card)


1
Expert's answer
2021-05-25T15:49:18-0400

"P(3 \,or \,red)=\frac{26+4-2}{52}=\frac{28}{52}=\frac{7}{13}."